OpenStudy (anonymous):

At what point do the equations y=x^2+6x+9 and y=-2x-3 intersect?

OpenStudy (anonymous):

set \[x^2+6x+9=-2x-3\] and solve for \(x\)

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

start with \[x^2+8x+12=0\] and then factor if you can or else use the quadratic formula yw

OpenStudy (anonymous):

So you would get (x+2)(x+6)=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the roots would be -6 and -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you just plug it back into the old equation to find y right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok I get it now. Thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

You're really helpful!

OpenStudy (anonymous):

thanks, glad to hear it